There were twelve times more children than adults at the matinee performance of the school play there were 455 people at the play. How many adults and children were there
Answer by fractalier(6550)

Call the number of children and adults, c and a.
Then we have
c = 12a and
c + a = 455
Substituting we have
12a + a = 455
13a = 455
a = 35 adults
c = 12(35) = 420 kids